# Learning Visual Hierarchies with Hyperbolic Embeddings

### v1

@@ Abstract
Structuring latent representations in a hierarchical manner enables models to learn patterns at multiple levels of abstraction. However, most prevalent image understanding models focus on visual similarity, and learning visual hierarchies is relatively unexplored. In this work, for the first time, we introduce a learning paradigm that can encode user-defined multi-level visual hierarchies in hyperbolic space without requiring explicit hierarchical labels. As a concrete example, first, we define a part-based image hierarchy using object-level annotations within and across images. Then, we introduce an approach to enforce the hierarchy using contrastive loss with pairwise entailment metrics. Finally, we discuss new evaluation metrics to effectively measure hierarchical image retrieval. Encoding these complex relationships ensures that the learned representations capture semantic and structural information that transcends mere visual similarity. Experiments in part-based image retrieval show significant improvements in hierarchical retrieval tasks, demonstrating the capability of our model in capturing visual hierarchies.
@@ 1 Introduction
Recent works have demonstrated the utility of hyperbolic representation spaces for capturing hierarchical relationships in an unsupervised setting [8, 37, 35]. This emergent latent structure is appealing; nevertheless, meaningful hierarchies are often task and data dependent, and aligning model behavior with such human-defined hierarchies is essential for many applications. To this end, we introduce a learning paradigm that can encode user-defined multi-level hierarchies using hyperbolic space. Specifically, we first define a general part-based image hierarchy using object and part level annotations within and across images. Then, we introduce a model capable of structuring the latent space to preserve the defined hierarchy using only image/object level information. To our knowledge, we are the first to encode user-defined visual hierarchies without relying on explicit hierarchical labels or additional modalities. Finally, we introduce an evaluation metric to effectively measure hierarchical image retrieval.
Note that, hierarchy is an asymmetric relationship and has a high branching factor (see Fig. 1). To this end, we adopt the hyperbolic geometry as it provides a continuous approximation of such tree-like structures [29, 40]. To enforce the hierarchy, we break it into pairwise entailment relationships between images, objects, and parts, at multiple levels within an image as well as across images at category level. For pairwise entailment (asymmetric), we adapt the recently proposed angle-based asymmetric distance metric [37] within the contrastive learning paradigm, and extend it to handle cases with multiple positive relationships. In contrast to symmetric distances such as the inner-product used in [30, 21, 13, 33], this angle-based distance offers an additional degree of freedom along the radial axis to form emergent structures in the latent space.
For experimentation, we build a dataset of visual hierarchies using the bounding box annotations of OpenImages [25]. Our dataset includes entailment relationships between scenes, objects, and parts, within a single image as well as across images at the category level. Similarly, for hierarchical retrieval evaluation, we use the full training set to create ground truth hierarchy trees per scene/object. Additionally, we design a metric for evaluating hierarchical retrieval based on the optimal transport distance between the label distribution of the retrieval set and ground truth label distribution within the hierarchy tree. This, along with precision and recall metrics, demonstrates that our method captures semantic and structural information transcending mere visual similarity. Furthermore, we show the superior generalization capabilities of our model via zero-shot image classification experiments.
@@
  • To our knowledge, for the first time, we introduce a new learning paradigm to effectively encode user-defined multi-level visual hierarchies in hyperbolic space that does not require explicit hierarchical labels.
@@
  • We demonstrate superior generalization capabilities of our model beyond the user-defined hierarchies via zero-shot image classification evaluation and ablation experiments.
@@ 2 Preliminaries
We briefly review essential concepts in hyperbolic geometry here. We refer the reader to  [38] for a comprehensive treatment. Hyperbolic spaces are Riemannian manifolds with constant negative curvature and are fundamentally different from Euclidean or spherical space which has zero or constant positive curvature, respectively. The negative curvature enables properties such as divergence of parallel lines and exponential volume growth towards the boundary [3]. This volume growth property makes hyperbolic space an ideal candidate for embedding hierarchical and graph structured data [29], and has found many machine learning applications.
@@
The Lorentz model is a way to represent a hyperbolic space. It frames the $d$-dimensional hyperbolic space $\mathbb{H}^{d}$ with curvature $c$ within an $(d+1)$-dimensional Euclidean space $\mathbb{R}^{d+1}$ as follows:
$\mathbb{H}^{d}=\left\{\mathbf{x}\in\mathbb{R}^{d+1}\mid\langle\mathbf{x},\mathbf{x}\rangle_{\mathbb{H}}=-1/c,x_{0}>0\right\}\ ,$ (1)
@@
$\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{H}}=-x_{0}y_{0}+\sum_{i=1}^{d}x_{i}y_{i}\ .$ (2)
@@
Geodesics.
Geodesics in hyperbolic space are the shortest paths between points, analogous to straight lines in Euclidean geometry. In the Lorentz model, geodesics are the intersections of planes through the origin with the hyperboloid.
@@
Figure 2: An example hierarchy tree. a) Image $I$ with object-level bounding boxes. Each bounding box is entailed by $I$. b) Hierarchy trees created via bounding box-to-bounding box entailment within $I$ (larger bounding boxes entail smaller ones). c) Cross-image hierarchy created by sampling $N$ bounding boxes (with corresponding object classes) from other images, which are then entailed by $I$.
@@ 3.1 Part-Based Image Hierarchy
Given an image dataset with bounding box and object class annotations, we define a part-based image hierarchy where the full scene images – typically containing multiple objects – represent the highest level in the hierarchy, while individual objects constitute progressively lower levels, entailed by the full image. In this, larger bounding boxes that significantly envelope smaller ones are considered to entail those smaller ones, establishing a nested hierarchy. In this way, from the full scene to the smallest bounding box, a hierarchy can be defined by recursively applying the entailment rule: if $B$ contained in $A$, then $A$ entails $B$, denoted as $A\to B$. An illustrative example is shown in Fig. 2, where an example hierarchy tree could be road scene $\to$ cyclist $\to$ bicycle $\to$ wheels.
@@
By applying this recursively, hierarchy trees can be formed as shown in Fig. 2b.
@@
As noted above, the part-based hierarchy is a tree structure and given an image (or a cropped bounding box), the tree can be traversed using the pairwise entailment relationships. For evaluation, we construct this hierarchy tree per scene/object on the full training set. This hierarchy tree would be used to evaluate the hierarchical image retrieval results. For instance, when considering parent-to-child retrieval, the tree would be traversed from the query image to its leaves and vice versa. More implementation details can be found in supplementary materials.
A subtle point to note is that, in Eq. 4, all the bounding boxes have the full image as a parent node. So if there are other enclosing bounding boxes, then a single bounding box can have multiple parent nodes. In this case, the parent node at the most granular level is chosen and other links are removed when constructing the tree. However, when training our model, the hierarchy tree is not used and only the pairwise entailment relationships are enforced. As we will clarify in the next section, our loss has the flexibility to arrange the embeddings to satisfy this tree structure without violating the pairwise entailment.
@@ 3.2 Angle-Based Entailment Loss
We require an asymmetric distance function to enforce pairwise entailment relationships in hyperbolic space. To this end, we adapt the recently proposed hyperbolic-angle-based entailment loss [37]. This is a smooth contrastive variant of the entailment cone loss [11]. It was originally proposed to align text and image modalities with flexibility to “accept” the modality gap along the radial axis. This flexibility is crucial even in our case, where we enforce tree-like hierarchical relationships only via pairwise entailment. The degree of freedom along the radial axis allows our model to structure embeddings in a hierarchical manner without violating the pairwise entailment.
Our loss is a bidirectional version of [37], as illustrated graphically in Fig. 3. In particular, given embeddings of an entailment pair, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, where $\mathbf{x}$ entails $\mathbf{y}$, we maximize the angles $\beta_{1}$ and $\alpha_{2}$. This enforces entailment in a bidirectional manner. The angles $\beta_{1}$ and $\alpha_{2}$ can be computed using the exterior angle as follows:
@@
$-\mathbb{E}_{\mathcal{D}}\!\left[\log\frac{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)}{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)+\sum_{\mathbf{y}^{-}\in\mathcal{N}_{i}}\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i}^{-})}{\tau}\right)}\right]\ ,$
where $\mathcal{N}_{i}$ denotes the set of samples that do not have an entailment relationship with the parent embedding $\mathbf{x}_{i}$. Here, $\kappa:\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$ is the similarity function, and $\tau$ is the temperature parameter. Now, our bidirectional entailment loss can be written as:
$L_{\text{angle}}(\mathcal{D})=L^{p\to c}(\mathcal{D},\beta_{1})+L^{c\to p}(\mathcal{D},\alpha_{2})\ ,$ (9)
@@
In our implementation, we use a shared image encoder for both parent and child embeddings. Following the reparametrization of [8], we encode the space component of the Lorentz model in the tangent space at origin, use the exponential map to project to the hyperboloid, and the time component is obtained via Eq. 3 to compute the loss.
@@
Figure 3: Learning multi-level hierarchies via contrastive entailment angle loss. Our model first encodes parent-to-child pairs into embeddings, then maximizes $\beta_{1}$ and $\alpha_{2}$ using our contrastive entailment angle loss.
@@ 3.3 Hierarchical Retrieval Evaluation
To evaluate retrieval performance on the hierarchy tree, we also introduce a metric that captures the label distribution in the dataset. This is important as different labels can have different numbers of instances and the standard metrics such as Precision@k and Recall@k are agnostic to it.
@@ 4 Related Work
Hyperbolic geometry allows for exponential volume expansion towards the boundary [3], making it effective for embedding hierarchical structures. For example, hierarchical trees can be embedded with low distortion in hyperbolic space using just two dimensions, which is not possible in Euclidean space even with an unbounded number of dimensions [40]. These advantages have led to significant research into leveraging hyperbolic representations for various data types, including molecular structures [48], 3D [4, 18, 45], images [21, 14, 13, 22, 35], text data [42, 42, 12, 49, 19], and vision-language data [8, 37, 23, 1].
Hyperbolic embeddings can be learned through standard deep learning layers [20] with hyperbolic projection [29] or using hyperbolic neural networks [12]. Many prior NLP, computer vision and knowledge graph studies learn hierarchies from partially order data [12, 43, 27], or minimizing geodesic distance or maximizing similarities [29, 30, 47, 13]. Ganea et al. [12] introduced an angle-based entailment cone loss which pushes child nodes into the cone emanating from the parent node embedding. This approach has been applied to both text [12] and image data [9] with label hierarchies. Recently, this hyperbolic entailment loss was adapted for contrastive learning to develop representations in vision-language models [8, 37, 1, 33]. However, such methods remain relatively unexplored in the image domain. We adapted the angle-based entailment loss from [37] to encode part-based image hierarchies. Unlike previous image models that depend on label hierarchies, such as from WordNet [9], hand labels [28] or simple scene-to-object hierarchies within the same image [13], we capture more complex multi-level part-based hierarchies directly from image data, going beyond visual similarity.
@@ 5 Experiments
In order to train and evaluate methods, we construct HierOpenImages, a novel dataset that contains pairwise part-based image hierarchies. It is built from the OpenImages dataset [25] following the steps in Sec. 3.1.
@@
We evaluate the performance of the proposed method on two popular visual encoders 1) CLIP ViT (B/16) [36], pretrained on large-scale image-text pairs from the Internet, and 2) MoCo-v2 (ResNet-50) [6], pretrained on ImageNet. The embeddings are projected to dimension 128 in the final layer. The hyperbolic model has a learnable curvature parameter. Each model is fine-tuned for a single epoch on HierOpenImages using the contrastive angle-based entailment loss. We denote the model fine-tuned in hyperbolic space as ‘Hyp’, in Euclidean space as ‘Euc’, and the models without fine-tuning as ‘ZeroShot’. Refer to supplementary materials for more experimental details and results.
5.1 Part-based Image Retrieval
Denoting the full image as parent and the bounding box as child, we evaluate retrieval tasks in both child-to-parent and parent-to-child directions. Taking child-to-parent as an example, for fine-tuned models, the retrieval is performed by finding the Top-K parent images corresponding to maximum $\alpha_{2}$ as denoted in Fig. 3. For zero-shot CLIP and MoCo-v2 models, we evaluate results based on maximal cosine similarity, which is consistent with their pre-training. We also evaluated the recent hyperbolic image model, namely HCL [13], on these tasks; however, it did not outperform its backbone zero-shot model, MoCo-v2, in the part-based retrieval task. Please see supplementary materials for more details.

### v2

# Learning Visual Hierarchies in Hyperbolic Space for Image Retrieval

@@ Abstract
Structuring latent representations in a hierarchical manner enables models to learn patterns at multiple levels of abstraction. However, most prevalent image understanding models focus on visual similarity, and learning visual hierarchies is relatively unexplored. In this work, for the first time, we introduce a learning paradigm that can encode user-defined multi-level complex visual hierarchies in hyperbolic space without requiring explicit hierarchical labels. As a concrete example, first, we define a part-based image hierarchy using object-level annotations within and across images. Then, we introduce an approach to enforce the hierarchy using contrastive loss with pairwise entailment metrics. Finally, we discuss new evaluation metrics to effectively measure hierarchical image retrieval. Encoding these complex relationships ensures that the learned representations capture semantic and structural information that transcends mere visual similarity. Experiments in part-based image retrieval show significant improvements in hierarchical retrieval tasks, demonstrating the capability of our model in capturing visual hierarchies.
@@ 1 Introduction
Recent works have demonstrated the utility of hyperbolic representation spaces for capturing hierarchical relationships in an unsupervised setting [9, 40, 38]. This emergent latent structure is appealing; nevertheless, meaningful hierarchies are often task and data dependent, and aligning model behavior with such human-defined hierarchies is essential for many applications. To this end, we introduce a learning paradigm that can encode multi-level hierarchies as entailment pairs in hyperbolic space. As a concrete example, we first define a general part-based image hierarchy using object and part level classification annotations within and across images. Then, we introduce a model capable of structuring the latent space to preserve the defined hierarchy using only image/object level information. To our knowledge, we are the first to encode multi-level complex visual hierarchies without relying on explicit hierarchical labels or additional modalities. Finally, we introduce an evaluation metric to effectively measure hierarchical image retrieval.
Note that, hierarchy is an asymmetric relationship and has a high branching factor (see Fig. 1). To this end, we adopt the hyperbolic geometry as it provides a continuous approximation of such tree-like structures [32, 43]. To enforce the hierarchy, we break it into pairwise entailment relationships between images, objects, and parts, at multiple levels within an image as well as across images at category level. For pairwise entailment (asymmetric), we adapt the recently proposed angle-based asymmetric distance metric [40] within the contrastive learning paradigm, and extend it to handle cases with multiple positive relationships. In contrast to symmetric distances such as the inner-product used in [33, 24, 14], this angle-based distance offers an additional degree of freedom along the radial axis to form emergent structures in the latent space.
For experimentation, we build a dataset of visual hierarchies using the bounding box annotations of OpenImages [28]. Our dataset includes entailment relationships between scenes, objects, and parts, within a single image as well as across images at the category level. Similarly, for hierarchical retrieval evaluation, we use the full training set to create ground truth hierarchy trees per scene/object. Additionally, we design a metric for evaluating hierarchical retrieval based on the optimal transport distance between the label distribution of the retrieval set and ground truth label distribution within the hierarchy tree. Combined with Recall@k metrics, this demonstrates that our method captures semantic and structural information, transcending mere visual similarity. Furthermore, to the best of our knowledge, our model is the first to generalize to out-of-domain image hierarchies, achieving strong performance on unseen and diverse datasets.
@@
  • To our knowledge, for the first time, we introduce a new learning paradigm to effectively encode user-defined multi-level complex visual hierarchies in hyperbolic space that does not require explicit hierarchical labels.
@@
  • We demonstrate superior generalization capabilities of our model beyond the user-defined hierarchies via out-of-domain unseen data evaluation and ablation experiments.
@@ 2 Preliminaries
We briefly review essential concepts in hyperbolic geometry here. We refer the reader to  [41] for a comprehensive treatment. Hyperbolic spaces are Riemannian manifolds with constant negative curvature and are fundamentally different from Euclidean or spherical space which has zero or constant positive curvature, respectively. The negative curvature enables properties such as divergence of parallel lines and exponential volume growth with radius [4]. This volume growth property makes hyperbolic space an ideal candidate for embedding hierarchical and graph structured data [32], and has found many machine learning applications.
@@
The Lorentz model is a way to represent a hyperbolic space. It embeds the $d$-dimensional hyperbolic space $\mathbb{H}^{d}$ with curvature $c$ within an $(d+1)$-dimensional Minkowski space $\mathbb{R}^{d,1}$, a pseudo-Euclidean space with one negative dimension, as follows:
$\mathbb{H}^{d}=\left\{\mathbf{x}\in\mathbb{R}^{d,1}\mid\langle\mathbf{x},\mathbf{x}\rangle_{\mathbb{H}}=-1/c,x_{0}>0\right\}\ ,$ (1)
@@
$\langle\mathbf{x},\mathbf{y}\rangle_{\mathbb{H}}=-x_{0}y_{0}+\sum_{i=1}^{d}x_{i}y_{i}\ .\vspace{-1ex}$ (2)
@@
Figure 2: An illustrative example image hierarchies. a) Image $I$ with object-level bounding boxes. Each bounding box is entailed by $I$. b) Hierarchies created via bounding box-to-bounding box entailment within $I$ (larger bounding boxes entail smaller ones). c) Cross-image hierarchy created by sampling $N$ bounding boxes with corresponding object classes from other images, which are then entailed by $I$. Find details of cross-image sampling in Sec. 3.1.
@@ 3.1 Part-Based Image Hierarchy
Given an image dataset with bounding box and object class annotations, we define a part-based image hierarchy where the full scene images – typically containing multiple objects – represent the highest level in the hierarchy, while individual objects constitute progressively lower levels, entailed by the full image. In this, larger bounding boxes that significantly envelope smaller ones are considered to entail those smaller ones, establishing a nested hierarchy. In this way, from the full scene to the smallest bounding box, a hierarchy can be defined by recursively applying the entailment rule: if $B$ contained in $A$, then $A$ entails $B$, denoted as $A\to B$. An illustrative example is shown in Fig. 2, where an example hierarchy could be road scene $\to$ cyclist $\to$ bicycle $\to$ wheels.
@@
By applying this recursively, a tree-like hierarchy can be formed as shown in Fig. 2b. Note that our model can encode any “user-defined hierarchy” represented as entailment pairs in Eq. (4). Part-based image hierarchy is one use case.
@@
As noted above, the part-based hierarchy forms a tree structure, where an image or cropped bounding box can be traversed using pairwise entailment relationships. For evaluating hierarchical image retrieval, we construct this hierarchy tree per scene/object automatically using the full training set. However, the model is trained solely on pairwise entailment relationships and does not use the hierarchy tree.
@@ 3.2 Angle-Based Entailment Loss
We require an asymmetric distance function to enforce pairwise entailment relationships in hyperbolic space. To this end, we adapt the recently proposed hyperbolic-angle-based entailment loss [40], a smooth contrastive variant of the entailment cone loss [12]. The angle-only loss, without distance constraints on image pairs, provides flexibility to be distributed along the radial axis, allowing embeddings to align with the tree structure while preserving pairwise entailment.
Our loss is a bidirectional version of [40], as illustrated graphically in Fig. 3. In particular, given embeddings of an entailment pair, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ in the tangent space, where $\mathbf{x}$ entails $\mathbf{y}$, we maximize the angles $\beta_{1}$ and $\alpha_{2}$. This enforces entailment in a bidirectional manner. The angles $\beta_{1}$ and $\alpha_{2}$ can be computed using the exterior angle as follows:
@@
$-\mathbb{E}_{\mathcal{D}}\!\left[\log\frac{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)}{\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i})}{\tau}\right)+\sum_{\mathbf{y}^{-}\in\mathcal{N}_{i}}\exp\left(\frac{\kappa(\mathbf{x}_{i},\mathbf{y}_{i}^{-})}{\tau}\right)}\right]$
where $\mathcal{N}_{i}$ denotes the set of samples that do not have an entailment relationship with the parent embedding $\mathbf{x}_{i}$. Here, $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the similarity function, and $\tau$ is a learnable temperature parameter initialized to 0.07 following [9]. Now, our bidirectional entailment loss can be written as:
$L_{\text{angle}}(\mathcal{D})=L^{p\to c}(\mathcal{D},\beta_{1})+L^{c\to p}(\mathcal{D},\alpha_{2})\ .$ (9)
@@
In our implementation, we use a shared image encoder for both parent and child embeddings. Following the reparametrization of [9], we encode the space component of the Lorentz model in the tangent space at origin and project it onto the hyperboloid using the exponential map, enabling contrastive entailment angle loss computation in hyperbolic space.
@@
Figure 3: Learning multi-level hierarchies via contrastive entailment angle loss. Our model first encodes parent-to-child pairs into embeddings with exponential mapping, then maximizes $\beta_{1}$ and $\alpha_{2}$ using our contrastive entailment angle loss in hyperbolic space.
@@ 3.3 Hierarchical Retrieval Evaluation
To evaluate retrieval performance on the hierarchy tree, we also introduce a metric that captures the label distribution in the dataset. This is important as different labels can have different numbers of instances and the standard metrics such as Recall@k is agnostic to it.
@@ 4 Related Work
Hyperbolic geometry allows for exponential volume growth with respect to the radius [4], making it effective for embedding hierarchical structures. This advantage has led to significant research into leveraging hyperbolic representations for various data types, including molecular structures [51], 3D [5, 21, 48], images [24, 16, 14, 25, 38], text data [45, 45, 13, 52, 22], and vision-language data [9, 40, 26, 2].
Hyperbolic embeddings can be learned through standard deep learning layers [23] with hyperbolic projection [32] or using hyperbolic neural networks [13]. Many prior NLP, computer vision and knowledge graph studies learn hierarchies from partially order data [13, 46, 30], or minimizing geodesic distance or maximizing similarities [32, 33, 50, 14]. Ganea et al. [13] introduced an angle-based entailment cone loss which pushes child nodes into the cone emanating from the parent node embedding. This approach has been applied to both text [13] and image data [10] with label hierarchies. Recently, this hyperbolic entailment loss was adapted for contrastive learning to develop representations in vision-language models [9, 40, 2, 36]. However, such methods remain relatively unexplored in the image domain. We adapted the angle-based entailment loss from [40] to encode part-based image hierarchies. Many previous works are limited to learning hierarchies for single-class images using predefined labeled hierarchies, such as ImageNet [31, 24] or hand-labeled data [10]. In this work, we propose a learning method that fine-tunes pre-trained models on large-scale datasets for general images (with multiple classes per image) without hierarchical labels. The most relevant approach, HCL [14], models simple scene-to-object hierarchies, whereas we capture more complex, multi-level part-based hierarchies directly from image data, extending beyond visual similarity.
@@ 5 Experiments
For training and evaluation, we construct HierOpenImages, a novel dataset containing pairwise part-based image hierarchies built from the OpenImages dataset [28]. We further evaluate generalization on out-of-domain unseen datasets and hierarchies on the LVIS [17] dataset and 10 popular single-class datasets.
@@
We evaluate the performance of learning multi-level image hierarchies using two popular visual encoders 1) CLIP ViT (B/16) [39], pretrained on large-scale image-text pairs from the Internet, and 2) MoCo-v2 (ResNet-50) [7], pretrained on ImageNet. Note that both CLIP and MoCo-v2 models are fine-tuned and evaluated in an image encoder only setting. We use these pretrained image models as our baseline and compare our proposed angle-based hyperbolic method against its Euclidean counterpart. Each model is fine-tuned for a single epoch on HierOpenImages using the proposed contrastive angle-based entailment loss, and $\dagger$ denotes fine-tuning. The retrieval distance function (Dist. Func.) aligns with the scoring function used during training. For example, minimizing hyperbolic angles (Hyp Ang.) for CLIP-hyp and Euclidean angles (Euc Ang.) for CLIP-euc model.
We also compare the state-of-the-art hyperbolic hierarchical image-only model HCL [14], which is trained on scene-to-object relationships from the OpenImages dataset [28] in both hyperbolic and Euclidean space. Accordingly, we evaluate the retrieval performance of HCL [14] using both hyperbolic distance and cosine similarity. For completeness, we also fine-tuned HCL [14] on the same HierOpenImages dataset and included it in our comparisons. Note that fine-tuning or evaluating with a text encoder is not possible on the current HierOpenImages dataset. Furthermore, previous image-only supervised methods, trained on predefined single-class image hierarchies [10, 24] are not directly comparable. These methods require all image classes to be predefined during training, which is not feasible for complex multi-object scenes in the OpenImages [28] and LVIS [17] datasets, where images contain multiple objects with diverse class labels.
5.1 Main Results